Given the two planes x-y+2z-1=0 and 3x+2y-6z+4=0. Find a parametric equation for the intersection. (4)

1
Expert's answer
2022-05-26T11:23:02-0400
"x-y+2z-1=0""3x+2y-6z+4=0"

Set "z=0," solve for "x" and "y"


"x-y-1=0""3x+2y+4=0"


"y=x-1""3x+2(x-1)+4=0"


"x=-0.4""y=-1.4"

Point "(-0.4, -1.4, 0)"


"\vec{n_1}\times\vec{n_2}=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 1& -1 & 2 \\
3 & 2 & -6
\end{vmatrix}"

"= \vec{i}\begin{vmatrix}
 -1 & 2 \\
 2 & -6
\end{vmatrix}- \vec{j}\begin{vmatrix}
 1 & 2 \\
 3 & -6
\end{vmatrix}+ \vec{k}\begin{vmatrix}
 1 & -1 \\
 3 & 2
\end{vmatrix}"

"=2 \vec{i}+12 \vec{j}+5 \vec{k}"

Parametric equation of a line is:


"x=-0.4+2t,\\ y=-1.4+12t, \\z=5t"